# Entropy-conservative high-order methods for high-enthalpy gas flows

### v1

@@ 3.3 Limiting case
This can be seen to be exactly the energy flux as given by Eqn. (75) in [36] (which in turn is based on the flux proposed by Chandrashekar [7]), up to a trivial change of variables from $T$ to $R_{\mathrm{specific}}T$.
This concludes the derivation of an algorithm for computing entropy-conserving flux functions for a gas with an arbitrary internal energy function.
@@ 3.4 Algorithm for pre-processing and flux computation
Here, an outline of the overall algorithm for the pre-computation of relevant tabulated values and subsquent estimation of the flux components is presented.
@@
$T_{min}>0$, $T_{max}>T_{min}$, $\Delta T>0$, $\varepsilon_{vibr}(T)$, $T(\varepsilon_{int})$
@@
$E_{min}\leftarrow\frac{1}{m}\left(\frac{5}{2}{kT_{min}}+\varepsilon_{vibr}(T_{min})\right)$triangleright Compute minimum internal energy
$E_{max}\leftarrow\frac{1}{m}\left(\frac{5}{2}{kT_{max}}+\varepsilon_{vibr}(T_{max})\right)$triangleright Compute maximum internal energy
$\Delta E\leftarrow(E_{max}-E_{min})/N$triangleright Compute uniform internal energy spacing
@@
     $E_{u,A}[i]\leftarrow E_{min}+i\Delta E$triangleright Set element of uniformly spaced energy array
@@
     $T_{\mathrm{inv},A}[i]\leftarrow T(E_{u,A}[i],T_{A}[i])$triangleright Compute temperature from energy with a non-linear solver using $T_{A}[i]$ as a starting guess
@@
         $T_{a}\leftarrow T_{A}[i-1]$
         $T_{b}\leftarrow T_{A}[i]$
         $T_{ab}\leftarrow\frac{1}{2}\left(T_{a}+T_{b}\right)$triangleright Compute endpoints for integration via Simpson’s rule
         $\eta_{A}[i]\leftarrow\eta_{A}[i-1]+\frac{\Delta T}{6}\left(c_{v}(T_{a})/T_{a}+4c_{v}(T_{ab})/T_{ab}+c_{v}(T_{b})/T_{b}\right)$triangleright Compute integral part of entropy using Simpson’s rule
@@
Algorithm 1 shows a pseudocode description of how the necessary tabulated quantities are precomputed. To distinguish between the flow variables arrays of tabulated values, we subscript the latter with $A$. All of the arrays $T_{A}$, $E_{A}$, $E_{u,A}$, $c_{v,A}$, $T_{\mathrm{inv},A}$, $\eta_{A}$ are of size $N+1$ and store the tabulated values. The user has to provide the minimum and maximum values of the temperature range, as well as the temperature step $\Delta T$. The functions for computation of the specific internal energy $\varepsilon_{\mathrm{int}}$ and the specific heat $c_{v,\mathrm{int}}$ from temperature are assumed to be known. It is usually not possible to compute the temperature from the internal energy, and thus, some non-linear solver needs to be used to populate the $T_{\mathrm{inv},A}$ array (by computing temperature from the uniformly discretized energy range $E_{u,A}$). While higher-order integration methods can be implemented instead of Simpson’s rule, the linear interpolation of specific heats and temperature (in dependence on energy) also places bounds on the minimum error achievable with the scheme. In general, it is assumed that the discretization step $\Delta T$ is chosen such that the error due to the approximations used is extremely small even using the proposed simple integration and interpolation rules.
@@
     $f_{e,i}\leftarrow(\varepsilon_{\mathrm{int},i}-E_{min})/\Delta E$
     $I_{e,i}\leftarrow\lfloor f_{e,i}\rfloor$triangleright Find position in array of energy
     $f_{e,i}\leftarrow f_{e,i}-I_{e,i}$triangleright Find linear interpolation coefficient
     $T_{i}\leftarrow(1-f_{e,i})\cdot T_{\mathrm{inv},A}[I_{e,i}]+f_{e,i}\cdot T_{\mathrm{inv},A}[I_{e,i}+1]$triangleright Find temperature from energy
@@
     $\eta_{i}\leftarrow(1-f_{T,i})\cdot\eta_{A}[I_{T,i}]+f_{T,i}\cdot\eta_{A}[I_{T,i}+1]$triangleright Find integral part of entropy from temperature
@@
Algorithm 2 shows the computation of the quantities required for the calculation of the fluxes as given by Eqns. (36)–(43). First, the temperature is computed from the internal energy, and is then used to approximate the specific heats and integral parts of the entropy of the left and right states. In order to avoid division by zero, in case the temperatures of the left and right states are close (as defined by some prescribed tolerance $r$, for example $r=0.5\Delta T$), the jumps in the integral part of the entropy and internal energy are simply replaced by their derivatives evaluated at $T=(T_{ll}+T_{rr})/2$. In a more general scenario (e.g., a multi-species flow), the linear interpolation of temperature from the energy would need to be replaced by a call to a non-linear solver.
3.5 Computational complexity and error analysis
We now analyze the computational cost of the algorithm and the error due to interpolation. Compared to the case of constant $c_{v}$, for each entropy-conservative flux computation, the following additional computations are required:
@@ 3.5 Computational complexity and error analysis
    computation of $T_{+}$, $T_{-}$,
@@
In the following analysis we disregard the cost of the computation of temperature from energy, as in the present work a single-species formulation is assumed. This allows for significantly faster computation as compared to using a non-linear solver, which is required in the more general multi-species case.
Computation of a linear interpolation of a function of some variable $x$ given $x$ requires the following floating-point operations: 1 subtraction, 1 division operation and 1 rounding operation (to get the index in the tabulated array), 1 subtraction operation (to get the relative position between the two closest array values), 2 multiplication operations, 1 subtraction and 1 summation operation (to perform the linear interpolation). In addition, 2 array access operations are needed. In total, 1 division, 1 round-off, 2 multiplications, 1 summation, 3 subtractions, and 2 array accesses are required. We need to perform linear interpolation twice: to compute $\eta(T_{+})$ and $\eta(T_{-})$. Whilst the pseudocode in Algorithm 2 contains 6 linear interpolations (temperature $T$, $c_{v}(T)$, and $\eta(T)$ for each of the left and right states), the specific heat is required only in case the temperatures of the left and right states are very close (in which case interpolation of the integral part of the entropy is not needed; vice versa, if $|T_{+}-T_{-}|>r$, the interpolation of the specific heats is not required. Therefore, with careful handling of the possible cases, no more than 4 linear interpolations are needed for any given pair of neighbouring flow states. Additionally, as already stated, the interpolation of temperature is neglected in the present analysis. Two additional subtractions are needed to compute the jumps in energy and the integral part of the entropy. Finally, 2 division operations are needed to compute ${c_{v}\left(T^{\ast}\right)}/{T^{\ast}}$, $c_{v}\left(T^{\ast\ast}\right)$.
So in total, we require 12 subtractions/additions, 4 multiplications, 2 round-offs, 4 divisions, and 4 array accesses. Two of the divisions can be re-formulated as multiplications, as the division is by a constant value of $\Delta T$ that is set at the start of the simulation. In that case, the number of floating-point operations is 12 subtractions/additions, 6 multiplications, 2 round-offs, 2 divisions.
@@
Having replaced division by multiplication where possible, the worst case cost of the flux evaluation is then 120 cycles, based on data on Ivy Bridge processor family latencies [10]. The formulation in [34] requires at least three exponentiations (each 68 cycles), two logarithms (each 90 cycles), 6 multiplications, 1 division, leading to an approximate cost of 424 cycles, approximately 3.5 as many cycles as the formulation in the present work. Again, it is important to note that in both cases, the cost of computing the temperature from the energy is neglected. Another factor not considered in the present analysis is the role of cache effects. The approach proposed in the present work requires storing several large one-dimensional arrays of floating point numbers, which are being accessed at every step of the simulation, which could lead to frequent loading and eviction of data from the CPU cache.
The sources of numerical error in the presented flux formulation are
  1. 1.
    Error in temperature from linear interpolation (computing temperature from energy)
  2. 2.
    Error in approximation of $\eta(T)$ when storing the tabulated values (dependent on quadrature rule used in the pre-processing step)
  3. 3.
    Error in the linear interpolation of the integral $\eta(T)$ between the tabulated values.
For the linear interpolation used in the present work, the error in the estimation of temperature is of second order in $\Delta E$. The error in the tabulated values of $\eta(T)$ is of fourth order in $\Delta T$ (assuming Simpson’s rule is used). Finally, the error in the linear interpolation of the integral part of the entropy is of second order in $\Delta T$. A numerical investigation of the errors in the flux evaluation (and the impact of different error sources), along with an analysis of the computational cost of the flux computation algorithm, is presented in the following section.
@@ 4 Numerical results
To verify the developed flux and apply it to simulation of two-dimensional high-speed flows, the developed algorithms for computation of flow properties and fluxes were implemented in Trixi.jl [40, 38], a modular framework for solving systems of hyperbolic equations using the DGSEM method. The strong stability preserving Runge-Kutta method SSPRK43 was used for time integration, as provided by the DifferentialEquations.jl library [35]. The simulation code is publicly available on Github [33].
@@ 4.1 Comparison to exact expressions for fluxes
Figure 2: Relative error in the energy flux for different temperature jumps as a function of the discretization step $\Delta T$, $\gamma=1.4$.
@@ 4.1.1 Calorically perfect gas
In the case of a calorically perfect gas, the error due to interpolation of temperature is exactly zero, as in that case temperature is a linear function of energy and the inversion is exact. Fig. 2 shows the error in the evaluation of the flux for the energy $F_{E}^{\mathrm{num},x}$ using Eqn. 39 computed via Algorithm 2 for various choices of $\Delta T$, as compared to the exact expression for the calorically perfect gas as given by Eqn. 45. In this case, $c_{v}=\frac{5}{2}\frac{k}{m}$ was assumed, corresponding to an adiabatic index $\gamma=1.4$. The densities and velocities of the left and right states were assumed to be equal ($\rho_{ll}=\rho_{rr}=3.8485\cdot 10^{-3}$ kg/m3, $v_{x,ll}=v_{x,rr}=1000$ m/s, $v_{y,ll}=v_{y,rr}=500.0$ m/s), the temperature of the left state $T_{ll}$ was taken to be 1000 K and four different temperature jumps were considered: $T_{rr}/T_{ll}=1.001$, $T_{rr}/T_{ll}=1.01$, $T_{rr}/T_{ll}=1.1$, and $T_{rr}/T_{ll}=2.0$. The tolerance $r$ used in Alg. 2 was taken to be equal to $\Delta T/2$; that is, for $\left|T_{rr}-T_{ll}\right|<r$ the specific heat was simply evaluated at the midpoint $\left(T_{rr}+T_{ll}\right)/2$, as opposed to using look-up tables for the integral part of the entropy and the internal energy. This is the reason why on Fig. 2 the error levels out after a certain $\Delta T$ — as soon as $\Delta T$ becomes large enough that $\left|T_{rr}-T_{ll}\right|<\Delta T/2$, the flux computation simply computes the specific heat at the mean temperature, and is thus no longer affected by the discretization error in the temperature. Since a constant $\gamma$ is used for this test case, there is no error due to the computation of temperature from the energy. It can be seen that for the case of a reasonable discretization step (i.e., less than 10 K), the error is on the order of 10-8–10-4 and decreases rapidly with smaller $\Delta T$.

### v2

@@ 3.3 Properties of the entropy-conservative flux
This can be seen to be exactly the energy flux as proposed by Chandrashekar [9] up to a trivial change of variables from $T$ to $R_{\mathrm{specific}}T$. ∎
Theorem 3.
Let $\varepsilon_{\mathrm{int}}(T)$ and $c_{v}(T)$ be piece-wise interpolated with error of order $(\Delta T)^{n}$, where $\Delta T$ is the spacing between equidistant values of $T$ used for the interpolation. We assume $\eta(T)$ to be computed exactly using the piece-wise interpolation for $c_{v}(T)$ (see Eqn. (46)). Let us also assume 1) a negligibly small error in the non-linear estimation of temperature from the interpolated energy, 2) that the function $\varepsilon_{\mathrm{int}}(T)$ is differentiable, 3) the derivative $c_{v}(T)=\varepsilon_{\mathrm{int}}^{\prime}(T)$ is bounded from below by a non-zero value. Then the the error in the numerical energy flux evaluated using Eqn. (41) using a piece-wise interpolation for $\varepsilon_{\mathrm{int}}(T)$ and $c_{v}(T)$ when compared to an exact computation of the flux (that is, it is assumed that one can compute $T(\varepsilon_{\mathrm{int}})$ and $\eta(T)$ with no numerical errors) is of the order $\left(\Delta T\right)^{n}$.
First, we prove the following lemma.
Lemma 1.
Let $p(x)$ be a piece-wise interpolation of a differentiable function $f(x)$ with error of order $(\Delta x)^{n}$, where $\Delta x$ is the spacing between equidistant values of $x$ used for the interpolation. Let $f^{\prime}(x)\geq A>0$, where $A$ is a constant. Then given a value $f_{0}=f(x_{0})$, the following holds: $\left|f^{-1}(f_{0})-p^{-1}(f_{0})\right|\leq C(\Delta x)^{n}$, where $C$ is a constant.
Proof.
Let us denote $p^{-1}(f_{0})$ as $\tilde{x}_{0}$. As $f$ is differentiable, we can apply the mean value theorem, assuming without loss of generality that $\tilde{x}_{0}>x_{0}$:
$\frac{f(\tilde{x}_{0})-f(x_{0})}{\tilde{x}_{0}-x_{0}}=f^{\prime}(c),\>c\in[x_{0},\tilde{x}_{0}].$ (53)
We have that $f(x_{0})=f_{0}=p(\tilde{x}_{0})$, so we can write that
$\left|\tilde{x}_{0}-x_{0}\right|=\frac{\left|f(\tilde{x}_{0})-f(x_{0})\right|}{\left|f^{\prime}(c)\right|}=\frac{\left|f(\tilde{x}_{0})-p(\tilde{x}_{0})\right|}{\left|f^{\prime}(c)\right|}\leq\frac{C_{1}(\Delta x)^{n}}{\left|f^{\prime}(c)\right|}\leq\frac{C_{1}}{A}(\Delta x)^{n}=C(\Delta x)^{n}.$ (54)
Now we prove the statement of the theorem.
Proof.
We consider the energy flux in the $x$ direction, as proof for the $y$ direction is done by analogy. Given two vector of conservative variables $\mathbf{u}_{-}$, $\mathbf{u}_{+}$, the values of $\rho$, $v_{x}$, $v_{y}$, and $\varepsilon_{\rm{int}}$ for the states -- and $+$ can be computed exactly.
Therefore the error comes only from the term $\left\{\!\left\{T\right\}\!\right\}_{\mathrm{geo}}^{2}\left(\frac{c_{v}\left(T^{\ast}\right)}{T^{\ast}}-\left\{\!\left\{\frac{1}{T}\right\}\!\right\}c_{v}\left(T^{\ast\ast}\right)\right)$.
Recalling the definitions of ${c_{v}\left(T^{\ast}\right)}/{T^{\ast}}$, $c_{v}\left(T^{\ast\ast}\right)$, we can re-write the above term as
$\left\{\!\left\{T\right\}\!\right\}_{\mathrm{geo}}^{2}\left(\frac{\eta(T_{+})-\eta(T_{-})}{T_{+}-T_{-}}-\left\{\!\left\{\frac{1}{T}\right\}\!\right\}\frac{\varepsilon_{\rm{int},+}-\varepsilon_{\rm{int},-}}{T_{+}-T_{-}}\right).$ (55)
According to Lemma 1, the error in the estimation of $T_{-}$ and $T_{+}$ is of order $(\Delta T)^{n}$. Therefore, the errors in the computation of $\left\{\!\left\{T\right\}\!\right\}_{\mathrm{geo}}^{2}$, $\left\{\!\left\{\frac{1}{T}\right\}\!\right\}$, and $\frac{1}{T_{+}-T_{-}}$ are also all of order $(\Delta T)^{n}$; this also holds for the error in the computation of $\varepsilon_{\rm{int},+}$, $\varepsilon_{\rm{int},-}$ and for any products of these quantities. Therefore, the error in $\left\{\!\left\{T\right\}\!\right\}_{\mathrm{geo}}^{2}\left\{\!\left\{\frac{1}{T}\right\}\!\right\}\frac{\varepsilon_{\rm{int},+}-\varepsilon_{\rm{int},-}}{T_{+}-T_{-}}$ is of order $(\Delta T)^{n}$.
The only remaining term to analyze is $\eta(T_{+})-\eta(T_{-})$. There are two sources of error: 1) the error due to the interpolation used for $c_{v}(T)$ and 2) the error due to the choice of interpolation points $T_{-}$, $T_{+}$ (as they are computed from an interpolation of energy). The errors due to the interpolation of $c_{v}(T)$ and the choice of integration limits are both of order $(\Delta T)^{n}$. Thus, the errors in the computation of $\eta(T_{+})$, $\eta(T_{-})$ are also of order $(\Delta T)^{n}$
Therefore, the term (55) appearing in the energy flux has error of order $(\Delta T)^{n}$. ∎
This concludes the analysis of the properties of the developed entropy-conserving numerical fluxes for a gas with an arbitrary internal energy function. To summarize:
  • The flux is consistent.
  • The flux is kinetic energy preserving.
  • The flux reduces to the entropy-conservative flux of Chandrashekar in case of a constant specific heat.
  • If piece-wise interpolation with a step-size of $\Delta T$ and error of order $(\Delta T)^{n}$ is used to compute the internal energy $\varepsilon_{\mathrm{int}}(T)$ and $c_{v}(T)$, whilst the integral part of the entropy $\eta$ is computed exactly using the piece-wise interpolation for $c_{v}(T)$, then the error in the energy flux is of the order $(\Delta T)^{n}$.
@@ 3.4 Algorithm for pre-processing and flux computation
Here, an outline of the overall algorithm for the pre-computation of relevant tabulated values and subsequent estimation of the flux components is presented.
@@
$T_{min}>0$, $T_{max}>T_{min}$, $\Delta T>0$, $\varepsilon_{\mathrm{int}}(T)$, $c_{v}(T)$
@@
         $\eta_{A}[i]\leftarrow\eta_{A}[i-1]+\left(c_{v,A}[i-1]-\frac{(c_{v,A}[i]-c_{v,A}[i-1])T_{A}[i]}{\Delta T}\right)\ln\left(\frac{T_{A}[i]}{T_{A}[i-1]}\right)+\left(c_{v,A}[i]-c_{v,A}[i-1]\right)$triangleright Compute integral part of entropy using piece-wise interpolation for $c_{v}$
@@
Algorithm 1 shows a pseudocode description of how the necessary tabulated quantities are precomputed. To distinguish between the flow variables arrays of tabulated values, we subscript the latter with $A$. All of the arrays $T_{A}$, $E_{A}$, $c_{v,A}$, $\eta_{A}$ are of size $N+1$ and store the tabulated values. The user has to provide the minimum and maximum values of the temperature range, as well as the temperature step $\Delta T$. The functions for computation of the specific internal energy $\varepsilon_{\mathrm{int}}$ and the specific heat $c_{v}$ from temperature are assumed to be known. Naturally, the discretization step $\Delta T$ should be chosen such that the error due to the approximations used is acceptably small; the impact of the choice of $\Delta T$ is discussed in the next section.
@@
     $T_{i}\leftarrow T(\varepsilon_{\mathrm{int},i})$triangleright Compute temperature from energy using a non-linear solver
@@
     $\eta_{i}\leftarrow\eta_{A}[I_{T,i}]+\left(c_{v,A}[I_{T,i}]-\frac{(c_{v,i}-c_{v,A}[I_{T,i}])T_{A}[I_{T,i}]}{\Delta T}\right)\ln\left(\frac{T}{T_{A}[I_{T,i}]}\right)+\left(c_{v,i}-c_{v,A}[I_{T,i}]\right)$triangleright Compute integral part of entropy
@@
Algorithm 2 shows the computation of the quantities required for the calculation of the fluxes as given by Eqns. (38)–(45). First, the temperature is computed from the internal energy, and is then used to approximate the specific heats and integral parts of the entropy of the left and right states. In order to avoid division by zero, in case the temperatures of the left and right states are close (as defined by some prescribed tolerance $r$, for example $r=0.5\Delta T$), the jumps in the integral part of the entropy and internal energy are simply replaced by their derivatives evaluated at $T=(T_{-}+T_{+})/2$. Such a choice of $T$ also ensures consistency of the flux in accordance with Theorem 1. It should be noted that the non-linear solver used to compute the temperature from energy also uses the the linear interpolation of the energy and specific heats in the iterations.
3.5 Computational complexity
We now briefly analyze the computational cost of the algorithm. Compared to the case of constant $c_{v}$, for each entropy-conservative flux computation, the following additional computations are required:
@@ 3.5 Computational complexity
    computation of $T_{+}$, $T_{-}$ from $\varepsilon_{\mathrm{int},+}$, $\varepsilon_{\mathrm{int},-}$ using a non-linear solver,
@@
In the following analysis we disregard the cost of the computation of temperature from energy (as this can be a call to an independent non-linear solver). For the case of $|T_{+}-T_{-}|<r$, only subtraction, division, and multiplication operations are required. For the case of $|T_{+}-T_{-}|\geq r$, additional operations are required, as two specific heats need to be linearly interpolated instead of one; in addition, two logarithms need to be computed. Some of the division operations, such as dividing by $\Delta T$, can be replaced by multiplication operations if the quantity $1/\Delta T$ is pre-computed and stored, leading to a lower computational effort [15]. The formulation in [42] requires at least three exponentiations and computing two logarithms, whereas the present algorithm does not require the exponentiation operations, thus leading to a potentially lower computational cost. Again, it is important to note that in this brief analysis, the cost of computing the temperature from the energy is neglected. In addition, it should again be stressed that the computational cost of the flux formulation proposed in the present work is independent of the exact expressions for the internal energy and specific heat used in the precomputation step. Another possible factor influencing the computational efficiency of the proposed algorithm not considered in the present analysis is the role of cache effects. The approach proposed in the present work requires storing several large one-dimensional arrays of floating point numbers, which are being accessed at every step of the simulation, which could lead to frequent loading and eviction of data from the CPU cache. Timing results of the actual implementations of the various flux functions are presented in the next section.
3.6 Dissipative flux
The entropy-conservative flux (38)–(45) can be augmented with a dissipation term and used to reduce the oscillations in flows with strong shocks. In principle, the dissipative term can be chosen such that flux is entropy-stable:
$\mathbf{F}^{\mathrm{num},ES,j}=\mathbf{F}^{\mathrm{num},EC,j}-\frac{1}{2}\mathbf{D}\llbracket\bm{\omega}\rrbracket,\>j=x,y.$ (56)
Here $\mathbf{D}$ is a dissipation matrix constructed in a way such that the flux $\mathbf{F}^{\mathrm{num},ES,j}$ is entropy-stable [13, 55]. The construction of a suitable dissipation operator $\mathbf{D}$ is left for future work. In the present work, for simulations with strong shocks a local Lax-Friedrichs scalar dissipation is added to the entropy-conservative flux:
$\mathbf{F}^{\mathrm{num},D,j}=\mathbf{F}^{\mathrm{num},EC,j}-\frac{\lambda}{2}\llbracket\mathbf{u}\rrbracket,\>j=x,y,$ (57)
where $\mathbf{F}^{\mathrm{num},ES,i}$ is the entropy-stable version of the entropy-conservative flux $\mathbf{F}^{\mathrm{num},EC,j}$, and $\lambda$ is the fastest local wave speed. This version of the flux is used for the simulations containing strong shocks presented in the next section.
@@ 4 Numerical results
To verify the developed flux and apply it to simulation of two-dimensional high-speed flows, the developed algorithms for computation of flow properties and fluxes were implemented in Trixi.jl [49, 45], a modular framework for solving systems of hyperbolic equations using the DGSEM method. The strong stability preserving Runge-Kutta method [18] SSPRK43 was used for time integration, as provided by the DifferentialEquations.jl library [43]. The simulation code is publicly available on Github [41].
@@ 4.1 Comparison to exact expressions for fluxes
Figure 2: Relative error in the energy flux for different temperature jumps as a function of the discretization step $\Delta T$ for a calorically perfect gas.
@@ 4.1.1 Calorically perfect gas
In the case of a calorically perfect gas, the error due to interpolation of temperature is almost zero, as in that case temperature is a linear function of energy and the inversion via the Newton-Solver is nearly exact. For this test case we assume that $c_{v}=\frac{5}{2}\frac{k}{m}$, which corresponds to an adiabatic index $\gamma=1.4$. The densities and velocities of the left and right states were assumed to be and given by $\rho_{-}=\rho_{+}=3.8485\cdot 10^{-3}$ kg/m3, $v_{x,-}=v_{x,+}=1000$ m/s, $v_{y,-}=v_{y,+}=500.0$ m/s, whereas the temperature of the left state $T_{-}$ was taken to be 1000 K and six different temperature jumps were considered: $T_{+}/T_{-}=1.001$, $T_{+}/T_{-}=1.01$, $T_{+}/T_{-}=1.1$, $T_{+}/T_{-}=2.0$, $T_{+}/T_{-}=5.0$, and $T_{+}/T_{-}=10.0$. The tolerance $r$ used in Alg. 2 was taken to be equal to $\Delta T/2$; that is, for $\left|T_{+}-T_{-}\right|<r$ the specific heat was simply evaluated at the midpoint $\left(T_{+}+T_{-}\right)/2$, as opposed to using look-up tables for the integral part of the entropy and the internal energy. In practice, significantly smaller values of the tolerance $r$ should be used, as it leads to a noticeable reduction in the error in the flux evaluation, as will be seen further in Sections 4.24.3. However, to better showcase how the flux behaviour changes once $\left|T_{+}-T_{-}\right|\geq r$, the relatively large value of $r=\Delta T/2$ was chosen for this test case.
Figure 2 shows the error in the evaluation of the flux for the energy $F_{E}^{\mathrm{num},x}$ using Eqn. 41 computed via Algorithm 2 for various choices of $\Delta T$, as compared to the exact expression for the calorically perfect gas as given by Eqn. 52, i.e. the entropy-conservative flux of Chandrashekar [9]. It can be observed that as long as $\left|T_{+}-T_{-}\right|\geq\Delta T/2$, the error in the flux computed via Algorithm 2 is extremely small, almost on the order of machine precision. It can also be seen that for a given temperature jump $T_{+}/T_{-}$, as soon as $\Delta T$ is chosen large enough that $\left|T_{+}-T_{-}\right|<\Delta T/2$, the error in the flux becomes significantly larger. Therefore, it can already be seen that using smaller values of the tolerance $r$ will lead to smaller errors, as in fewer cases will the simplified computation be used. For reasonable choices of a temperature discretization step (i.e. 1–10 K), the error still remains low even when the flux is computed using the simplified approximation with the mean temperature.